Solve by the addition method. (If there is no solution, enter NO SOLUTION. Use the parameters x and y as necessary.)
3x/4 - 2/y = 0
5x/4 - y/3 = 7/36

Answer by ikleyn(53763)

.
Solve by the addition method. (If there is no solution, enter NO SOLUTION. Use the parameters x and y as necessary.)
3x/4 - 2/y = 0
5x/4 - y/3 = 7/36
~~~~~~~~~~~~~


Double check and re-check your post.

The system does not seem to be written correctly.